# NeuralOM: Neural Ocean Model for Subseasonal-to-Seasonal Simulation

### v2

@@ 3.3 Optimization
We apply the relative $L_{2}$ loss function to train each component of NeuralOM, which can be defined as:
$\mathcal{L}_{2}=\frac{1}{KHW}\sum_{k=1}^{K}\sum_{i=1}^{H}\sum_{j=1}^{W}\frac{\left(\hat{O}_{i,j,k}^{t+l\delta t}-O_{i,j,k}^{t+l\delta t}\right)^{2}}{\left(O_{i,j,k}^{t+l\delta t}\right)^{2}},$ (29)
where, $\hat{O}_{i,j,k}^{t+l\delta t}$ and $O_{i,j,k}^{t+l\delta t}$ are the predicted and ground truth for ocean variables (channels) $k$ at spatial location $(i,j)$ and time $t+l\delta t$; $K$ is the number of variables; $H$ and $W$ are the height and width of the spatial dimensions, respectively; $\delta t$ is the time interval of single-step simulation (we use $\delta t=24$ hours).
3.4 Global Ocean Simulation
As illustrated in Figure 1(c), our inference procedure follows a rollout strategy: starting from the initial state $\mathbf{X}_{0}$, the trained model autoregressively feeds its own previous simulations together with the corresponding external forcing into the next time step. Previous studies have shown that evaluating an ocean model with forecast atmospheric fields can blur the distinctions among different ocean models, because errors in the atmospheric forecasts tend to dominate and “smooth out” their performance differences wang2024xihe  shu2025advanced . In other words, when an ocean-only model is driven by biased forecast forcing, a well-designed ocean model can appear worse, while a poorer one can appear better. Consequently, a more reliable assessment of an ocean model’s skill is obtained by using ground truth atmospheric forcing rather than forecast fields. Therefore, we use the ground truth atmospheric forcing to drive ocean models for comprehensive comparisons.
4 Experiment
In this section, we thoroughly evaluate the performance of NeuralOM, including metric results and visual results. We conduct all experiments on 64 NVIDIA A100 GPUs.
Table 1: In the global ocean simulation task, we compare the performance of our NeuralOM with 3 baselines. The average results for all 92 ocean variables of RMSE and ACC are recorded. Given the magnitude of the values for different variables, we report the normalized RMSE. A smaller RMSE ($\downarrow$) and a bigger ACC ($\uparrow$) indicate better performance. The best results are in bold, and the second best are with underline.
Model Metric
@@ 4 Experiment
FourCastNet pathak2022FourCastNet arXiv 2022 5.6336 0.0781 >10 0.0189 >10 0.0046 >10 0.0007
CirT liu2025cirt ICLR 2025 2.7525 0.0143 2.8700 0.0115 2.9528 0.0099 3.0645 0.0081
WenHai cui2025forecasting Nature Communications 2025 0.7200 0.5067 0.7468 0.4705 0.7702 0.4386 0.8093 0.3852
NeuralOM (ours) 0.6619 0.5210 0.6785 0.4916 0.6926 0.4656 0.7158 0.4212
NeuralOM (Promotion) 8.07% 2.84% 9.15% 4.49% 10.07% 6.16% 11.55% 9.33%
@@
Figure 3: The latitude-weighted RMSE (lower is better) and ACC (higher is better) results of several important ocean surface variables.
@@
4.1 Benchmarks and Baselines
We conduct the experiments on GlORYS12 reanalysis data, which offers daily mean data covering latitudes between -80° and 90° spanning between 1993 and 2023. The subset we use includes years from 1993 to 2020, which is 1993-2017 for training, 2018-2019 for validating, and 2020 for testing. We use 4 depth level ocean variables (each with 23 depth levels, corresponding to 0.49 m, 2.65 m, 5.08 m, 7.93 m, 11.41 m, 15.81 m, 21.60 m, 29.44 m, 40.34 m, 55.76 m, 77.85 m, 92.32 m, 109.73 m, 130.67 m, 155.85 m, 186.13 m, 222.48 m, 266.04 m, 318.13 m, 380.21 m, 453.94 m, 541.09 m and 643.57 m), Sea salintiy (S), Sea stream zonal velocity (U), Sea stream meridonal velocity (V), Sea temperature (T), and 1 surface level variable Sea surface height (SSH). We downsampling them to 1/2 degree (H=361, W=720). And we use the data with size 360 × 720. 4 atmosphere variables are used as forcing. More details can be found in the Appendix. We compare the simulation performance of NeuralOM with 3 data-driven models (FourCastNet pathak2022FourCastNet , CirT liu2025cirt , and WenHai cui2025forecasting ). FourCastNet is an atmospheric backbone model with good performance. CirT is a SOTA subseasonal-to-seasonal atmosphere prediction model that is similar to our simulation time scale. WenHai is a SOTA ML-based ocean model that outperforms traditional numerical models and ML-based models. Specifically, we directly use the inference code released by WenHai’s author to produce the simulation results of WenHai. For other baselines and our NeuralOM, we retrain them in the same framework and settings.
4.2 Evaluation Metrics
where $\mathbf{A}_{i,j,t}^{\mathcal{K}}$ represents the value of variable $\mathcal{J}$ at horizontal coordinate $(i,j)$ and time t. Latitude-dependent weights are defined as $L(i)=N_{\text{lat }}\times\frac{\cos\phi_{i}}{\sum_{i\textquoteright=1}^{N_{\text{lat}}}\cos\phi_{i\textquoteright}}$, where $\phi_{i}$ is the latitude at index i. The anomaly of $A$, denoted as $A^{\prime}$, is computed as the deviation from its climatology, which corresponds to the long-term mean of the meteorological state estimated from multiple years of training data. RMSE and ACC are averaged across all time steps and spatial coordinates, providing summary statistics for each variable $\mathcal{J}$ at a given lead time $\Delta t$.
where, true positives (TP) represent the number of instances in which the ocean state is correctly simulated, while false positives (FP) and false negatives (FN) are defined analogously. $F=\frac{\mathrm{FP}}{\mathrm{FP}+\mathrm{TP}}$ is the false alarm rate, and $H=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}}$ represents the hit rate.
4.3 Global Ocean Simulation
Given the magnitude of the values for different variables, we report the normalized RMSE. We first normalize the 93 ocean variables and then calculate the indicators for the 240 initial conditions (ICs) starting from 1 Jan. 2020, and each IC will be updated one day later.
Figure 4: The CSI and SEDI results of extreme event assessment.
As shown in Table 1, NeuralOM achieves competitive performance compared with the state-of-the-art models. As shown in Figure 2, NeuralOM is closer to the ground truth. We also show the simulation results of several important surface variables in Figure 3, which are not normalized. It can be seen that state-of-the-art atmosphere prediction model, such as CirT, perform poorly in ocean scenes. A main reason may lays on the different nature of atmosphere and ocean dynamic systems. In contrast, our proposed NeuralOM performs satisfactorily in subseasonal-to-seasonal ocean simulation. This improvement is primarily attributed to integrating the proposed ocean-specific multi-stage framework and multi-scale interactive graph neural network, which effectively captures the ocean’s small change nature and accurately models the evolution of ocean dynamics at different time scales. More results at other depth-level layers can be found in the Appendix. It is worth noting that our NeuralOM achieves competitive results compared with the SOTA WenHai model training using 1/12 degree data, although we train NeuralOM using lower resolution data with 1/2 degree.
4.4 Extreme Events Assessment
We apply CSI and SEDI to evaluate the simulation ability for extreme events. As shown in Figure 4, we calculate the CSI and SEDI results of U0 and V0 for 50 ICs. A bigger value means better performance to extreme event simulation. Our proposed NeuralOM achieves competitive performance.
4.5 Ablation Studies
To assess the effectiveness of the proposed method, we conduct detailed ablation studies, as summarized in Table 2 and Table 3. Specially, we introduce several model variants: (1) NeuralOM W/O Subtract Climatology, we don’t subtract climatology for periodicity variables before input them to the model.
Variants RMSE ACC
NeuralOM w/o OMF 0.6692 0.5503
NeuralOM w/o PEM 0.6865 0.5168
NeuralOM w/o MNM Sum 0.7037 0.5361
NeuralOM w/o MNM Mean 0.6978 0.5161
NeuralOM 0.6097 0.5983
(2) NeuralOM w Subtract Climatology, we subtract the climatology for the periodicity variables. (3) NeuralOM w/o OMF, we remove ocean-specific multi-stage framework. (4) NeuralOM w/o PEM, we remove physical process-based edge update module. (5) NeuralOM w/o MNM Sum, we remove the sum aggregation strategy in the multi-scale node update module. (6) NeuralOM w/o MNM Mean, we remove the mean aggregation strategy in the multi-scale node update module. (7) NeuralOM, the full version of NeuralOM. The above metrics are based on 30-day simulation results (50ICs for (1),(2), and 100ICs for others). Experimental results show that removing any single component impairs NeuralOM’s performance, thereby validating the effectiveness of our proposed method.
5 Conclusion
This work presents NeuralOM, a neural ocean foundation model tailored for Subseasonal-to-Seasonal (S2S) simulation. Our proposed ocean-specific multi-stage framework ensures accurate modeling of the ocean’s slow dynamics, while the multi-scale interactive messaging module enables the representation of intricate dynamical interactions, which accurately model the evolution process of ocean dynamics at the S2S time scale. Extensive experimental results demonstrate that NeuralOM significantly outperforms existing state-of-the-art models in both S2S and extreme event simulation, confirming its potential to revolutionize ocean and climate simulation tasks.
@@ A.1 Dataset
We conduct the experiments on GlORYS12 reanalysis data, offering daily mean data covering latitudes between -80° and 90° spanning between 1993 and 2023, which can be downloaded from https://data.marine.copernicus.eu/product/GLOBAL_MULTIYEAR_PHY_001_030/description. The subset we use includes years from 1993 to 2020, which is 1993-2017 for training, 2018-2019 for validating, and 2020 for testing. We use 4 depth level ocean variables (each with 23 depth levels, corresponding to 0.49 m, 2.65 m, 5.08 m, 7.93 m, 11.41 m, 15.81 m, 21.60 m, 29.44 m, 40.34 m, 55.76 m, 77.85 m, 92.32 m, 109.73 m, 130.67 m, 155.85 m, 186.13 m, 222.48 m, 266.04 m, 318.13 m, 380.21 m, 453.94 m, 541.09 m and 643.57 m), Sea salinity (S), Sea stream zonal velocity (U), Sea stream meridional velocity (V), Sea temperature (T), and 1 surface level variable Sea surface height (SSH). To improve computational efficiency, we use bilinear interpolation to downsampling them to 1/2 degree (H=361, W=720). And to better adapt to the input of different architecture models, we use the data with size 360 × 720. 4 atmosphere variables from ERA5 are used as forcing, which include 10 metre u wind component (U10M), 10 metre v wind component (V10M), 2 metre temperature (T2M), and mean sea level pressure (MSLP). The ERA5 data can be downloaded from https://cds.climate.copernicus.eu/, the official website of Climate Data Store (CDS). All the data we used are shown in Table 4.
@@ A.2 Data preprocessing
Appendix B Algorithm
We summarize the overall framework of the ocean-specific multi-stage framework (OMF) in Algorithm 1 and multi-scale interactive graph neural networks (MIGNN) in Algorithm 2.
Algorithm 1 Ocean-specific Multi-stage Framework (OMF)
0:  Initial input data $X_{t}$.
0:  Next step ocean state $O_{t+1}$.
1:  Initialize OMF
2:  repeat
3:     MIGNN (1)
4:     Apply the first MIGNN model to encode the combined input $X_{t}=[O_{t},F_{t},F_{t+1}]$ into the intermediate results of the next step ocean state $\hat{O}^{1}_{t+1}={\mathcal{F}^{M}_{1}}(X_{t})$.
5:     MIGNN (2)
6:     Apply the second MIGNN model to map $\hat{O}^{1}_{t+1}$ into the intermediate results: $\hat{O}^{2}_{t+1}={\mathcal{F}^{N}_{2}}(\hat{O}^{1}_{t+1})+\hat{O}^{1}_{t+1}$.
7:     MIGNN (Q)
8:     Apply the Q-th MIGNN model to map $\hat{O}^{Q-1}_{t+1}$ into the final results: ${O}_{t+1}={\mathcal{F}^{N}_{Q}}(\hat{O}^{Q-1}_{t+1})+\hat{O}^{Q-1}_{t+1}$.
9:  until converged
10:  return $OMF$
Algorithm 2 Multi-scale Interactive Graph Neural Networks (MIGNN)
0:  Initial input data $X_{t}$.
0:  Next step ocean state $O_{t+1}$.
1:  Initialize MIGNN
2:  repeat
3:     Encoder
4:     Embedding features of grid nodes $X_{t}$, mesh nodes $\mathcal{V}$, mesh edges $\mathcal{E}_{m}$, grid to mesh edges $\mathcal{E}^{\mathrm{G}2\mathrm{M}}$, and mesh to grid edges $\mathcal{E}^{\mathrm{M}2\mathrm{G}}$ into latent space using separate MLP: ($\mathcal{V}_{f}^{G}$, $h$, $\mathcal{E}_{f}$, $\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}$, $\mathcal{E}^{\mathrm{M}2\mathrm{G}}_{f}$) = MLPs($X_{t}$, $\mathcal{V}$, $\mathcal{E}_{m}$, $\mathcal{E}^{\mathrm{G}2\mathrm{M}}$, $\mathcal{E}^{\mathrm{M}2\mathrm{G}}$)
5:     Project the ocean state from the lat-lon grid into the mesh nodes: ${\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}^{\prime}=\mathrm{ESMLP}(\mathcal{V}_{f}^{G},h,\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f})$, $h^{\prime}=\mathrm{MLP}_{e1}(h,\sum{\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}^{\prime})$
6:     Update grid node feature: ${\mathcal{V}_{f}^{G}}^{\prime}=\mathrm{MLP}_{e2}(\mathcal{V}_{f}^{G})$
7:     Apply residual connection to update the feature of grid to mesh edge, mesh node, and grid node again: ${\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}={\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}+{\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}^{\prime}$, $h^{\prime}=h+h^{\prime}$, ${\mathcal{V}_{f}^{G}}={\mathcal{V}_{f}^{G}}+{\mathcal{V}_{f}^{G}}^{\prime}$
8:     Multi-scale Interactive Massaging (MIM)
9:     Apply physical process-based edge update module (PEM) to update edge feature: $\mathcal{E}_{f}^{\prime}=PEM(\mathcal{E}_{f},h_{s},h_{r})$
10:     Apply multi-scale node update module (MNM) to update mesh node feature: $h^{\prime}=MNM(h,\sum{\mathcal{E}_{f}}^{\prime})$
11:     Apply residual connection to update the feature of edge and mesh node: $\mathcal{E}_{f}=\mathcal{E}_{f}+\mathcal{E}_{f}^{\prime}$, $h=h+h^{\prime}$
12:     Decoder Project the feature from mesh back to lat-lon grid: ${\mathcal{E}^{\mathrm{M}2\mathrm{G}}_{f}}^{\prime}=\mathrm{ESMLP}(\mathcal{V}_{f}^{G},\mathcal{E}_{f},\mathcal{E}^{\mathrm{M}2\mathrm{G}}_{f})$, ${\mathcal{V}_{f}^{G}}^{\prime}={\mathrm{MLP}_{d1}}(\mathcal{V}_{f}^{G},\sum{\mathcal{E}^{\mathrm{M}2\mathrm{G}}_{f}}^{\prime})$, $\mathcal{V}_{f}^{G}=\mathcal{V}_{f}^{G}+\mathcal{V}_{f}^{G^{\prime}}$, $O_{t+1}=\mathrm{MLP_{d2}}({\mathcal{V}_{f}^{G}})$
13:  until converged
14:  return $MIGNN$
Appendix C Model Training

### v3

@@ Introduction
We train each component of NeuralOM by minimizing the relative mean squared error (MSE) between the predicted and ground-truth states. The loss function $\mathcal{L}$ is defined as:
$\mathcal{L}(\Theta)=\mathbb{E}_{t,i,j,k}\left[\frac{\left(\hat{O}_{i,j,k}^{t+1}-O_{i,j,k}^{t+1}\right)^{2}}{\mathbb{E}_{i,j,k}[(O_{i,j,k}^{t+1})^{2}]+\epsilon}\right],$ (6)
where $\hat{O}$ and $O$ are the predicted and ground-truth values, respectively, indexed over time $t$, spatial locations $(i,j)$, and variable channels $k$. The expectation in the denominator serves as a normalization constant to balance the influence of variables with different magnitudes, and $\epsilon$ is a small constant for numerical stability.
Inference Strategy.
For long-range simulation and forecasting, we employ a standard autoregressive (or ”rollout”) strategy. Starting from an initial state $\mathbf{X}_{0}$, the model continuously feeds its own previous prediction as input for the next step to generate a continuous trajectory spanning several weeks or months. For idealized simulation experiments, we use ground-truth atmospheric data as forcing. This approach eliminates interference from atmospheric forecast errors, allowing for a fairer assessment of the ocean models’ intrinsic performance. For more realistic forecasting experiments, we drive NeuralOM with forecast fields generated by a state-of-the-art atmospheric forecasting model, OneForecast (Gao et al. 2025).
Summary.
In conclusion, the proposed NeuralOM framework forms an end-to-end solution for slow-changing physical systems, bridging macroscopic correction strategies with microscopic physical modeling. To transparently illustrate its computational flow, we outline the complete forward pass of NeuralOM in Algorithm 1. In the following section, we will show its practical effectiveness through a series of rigorous benchmarks and ablation studies.
Algorithm 1 NeuralOM Forward Pass
1:Initial state $\mathbf{O}_{0}$, forcing sequence $\{\mathbf{F}_{t}\}_{t=0}^{T}$, total steps $T+1$
2:Predicted trajectory $\{\hat{\mathbf{O}}_{t}\}_{t=1}^{T}$
[…]
4:for $t=0,1,\dots,T-1$ do
5:  $\mathbf{X}_{t}\leftarrow\text{Concat}(\mathbf{O}^{\text{prev}},\mathbf{F}_{t},\mathbf{F}_{t+1})$ $\triangleright$ Combine current inputs
6: # Stage 1: Progressive Residual Correction Framework
7:  $\hat{\mathbf{O}}_{t+1}^{(0)}\leftarrow\text{PGN}_{\text{base}}(\mathbf{X}_{t})$ $\triangleright$ Initial prediction by the base model
8:  $\hat{\mathbf{O}}_{\text{intermediate}}\leftarrow\hat{\mathbf{O}}_{t+1}^{(0)}$
9:  for $q=1,\dots,Q$ do $\triangleright$ Perform Q rounds of residual correction
10:   $\text{Residual}\leftarrow\text{PGN}_{\text{residual}}^{(q)}(\hat{\mathbf{O}}_{\text{intermediate}})$ $\triangleright$ Residual model predicts the correction
11:   $\hat{\mathbf{O}}_{\text{intermediate}}\leftarrow\hat{\mathbf{O}}_{\text{intermediate}}+\text{Residual}$ $\triangleright$ Apply the residual correction
12:  end for
13:  $\hat{\mathbf{O}}_{t+1}\leftarrow\hat{\mathbf{O}}_{\text{intermediate}}$ $\triangleright$ Get the final single-step prediction
14:  $\mathbf{O}^{\text{prev}}\leftarrow\hat{\mathbf{O}}_{t+1}$ $\triangleright$ Update state for the next autoregressive step
15:end for
16:return $\{\hat{\mathbf{O}}_{t}\}_{t=1}^{T}$
Model Simulate Horizon
@@
FourCastNet (Pathak et al. 2022) arXiv 2022 5.6336 0.0781 >10>10 0.0189 >10>10 0.0046 >10>10 0.0007
CirT (Liu et al. 2025) ICLR 2025 2.7525 0.0143 2.8700 0.0115 2.9528 0.0099 3.0645 0.0081
WenHai (Cui et al. 2025) Nature Communications 2025 0.7200 0.5067 0.7468 0.4705 0.7702 0.4386 0.8093 0.3852
NeuralOM (ours) 0.6510 0.5299 0.6666 0.5010 0.6797 0.4756 0.7015 0.4316
NeuralOM Improvement 9.59% 4.58% 10.74% 6.49% 11.74% 8.43% 13.31% 12.02%
Table 1: Performance comparison of NeuralOM against baselines on global ocean simulation task. Average RMSE (normalized, $\downarrow$) and ACC ($\uparrow$) across all 93 ocean variables are shown. Best results in bold, second best underlined.
@@
Figure 4: The latitude-weighted RMSE and ACC results of several important ocean surface variables.
Experiments
To comprehensively evaluate the performance of NeuralOM and validate its effectiveness in addressing the challenges of long-term stability and physical consistency, we design a series of rigorous experiments. We first benchmark NeuralOM against several state-of-the-art models on an idealized global ocean simulation task. We then apply it to a more challenging real-world forecasting scenario. Finally, we dissect the contributions of our model’s innovative components through extensive ablation studies. All experiments were conducted on 64 NVIDIA A100 GPUs.
Experimental Setup
Dataset and Task.
Our experiments are based on the GLORYS12 global ocean reanalysis dataset, which provides daily mean ocean states from 1993 to 2023. We split the data into a training set (1993-2017), a validation set (2018-2019), and a test set (2020). The core task is autoregressive forecasting at the Subseasonal-to-Seasonal (S2S) scale, with rollout horizons ranging from several days up to 60 days. The input consists of 93 ocean state variables (including salinity, velocity, temperature at various depths, and sea surface height) and 8 atmospheric variables as external forcing. All spatial data is downsampled to a 0.5° resolution. Further details on data processing and variables are available in the Appendix.
Baselines.
We compare NeuralOM against three representative models: (1) FourCastNet (Pathak et al. 2022), a widely recognized atmospheric science foundation model based on Fourier Neural Operators. (2) CirT (Liu et al. 2025), a state-of-the-art atmospheric forecasting model designed for the S2S scale. (3) WenHai (Cui et al. 2025), a state-of-the-art ML-based ocean model that has demonstrated superior performance over traditional numerical models and ML-based ocean foundation models. To ensure a fair comparison, all models, except for the officially released WenHai results, were retrained under our unified framework.
Evaluation Metrics.
We use two primary metrics to assess performance: (1) Root Mean Square Error (RMSE) to measure the average prediction error, and (2) Anomaly Correlation Coefficient (ACC) to evaluate the model’s ability to capture deviations from the climatological mean (i.e., anomalies). We report the normalized RMSE and ACC, averaged over all 93 ocean variables. Additionally, we use the Critical Success Index (CSI) and the Symmetric Extremal Dependence Index (SEDI) to evaluate the model’s performance on extreme events.
Main Results and Analysis
NeuralOM sets a state-of-the-art in long-term simulation.
We first evaluate NeuralOM in an idealized simulation setting using ground-truth atmospheric forcing. As shown in Table 1, NeuralOM consistently outperforms all baselines. At a 60-day lead time, it reduces RMSE by 13.313.3% and improves ACC by 12.012.0% over the strongest baseline, WenHai. This superior performance is visually confirmed in Figure 3, where NeuralOM is the only model to maintain physical consistency and avoid the error cascades that cause other models to collapse. The significantly lower error growth rate, detailed in Figure 4, further underscores its long-term stability. The poor performance of powerful atmospheric models like CirT ($60$-day ACC of $0.0081$) also highlights the unique challenges of ocean simulation and validates our specialized design.
Robust performance is maintained in ocean forecasting.
To assess practical applicability, we drive the ocean models with atmospheric forecasts from OneForecast, which introduces real-world uncertainty. As common atmospheric forecast models don’t provide all forcings required by WenHai at the S2S scale, we can’t assess its results. Even in this more challenging scenario, Table 2 shows that NeuralOM retains its lead, achieving significantly better RMSE ($0.68$) and ACC ($0.49$) at a $35$-day lead time. This demonstrates the robustness of NeuralOM for real-world applications, further supported by the high-fidelity visuals in Figure 5(a).
Models Metrics
15-day 25-day 35-day
RMSE ACC RMSE ACC RMSE ACC
FourCastNet 0.56 0.68 0.69 0.47 1.54 0.16
CirT 1.61 0.11 2.17 0.04 2.55 0.02
Ours 0.53 0.73 0.62 0.60 0.68 0.49
Table 2: In the real-world global ocean forecasting task, we compare the performance of our NeuralOM with 2 baselines. The average results for all 93 ocean variables of RMSE and ACC are recorded. The best results are in bold.
Variants RMSE ACC
NeuralOM w/o PRC 0.6692 0.5503
NeuralOM w/o PEI 0.6865 0.5168
NeuralOM w/o MANA Sum 0.7037 0.5361
NeuralOM w/o MANA Mean 0.6978 0.5161
NeuralOM 0.6097 0.5983
NeuralOM excels at capturing high-impact extreme events.
Beyond average metrics, a model’s skill in predicting extreme events is critical. We use CSI and SEDI to evaluate predictions of extreme surface currents. As shown in Figure 5(b), NeuralOM achieves higher scores than all baselines at both 30 and 60-day lead times. This enhanced capability to capture potentially hazardous dynamics demonstrates its value for critical applications like risk assessment.
Figure 5: Performance on forecasting and extreme event assessment. (a) Visualizations of 30-day forecasts. Our model’s outputs align closely with the ground truth, while baseline models exhibit significant artifacts. (b) CSI and SEDI scores for extreme surface current events. NeuralOM achieves higher scores than baselines at both 30 and 60-day lead times.
@@
Ablation Studies
To deeply understand the driving factors behind NeuralOM’s superior performance and to validate the effectiveness of each of our design choices, we conduct a series of detailed ablation studies. We primarily investigate two aspects: the importance of the input representation (i.e., subtracting climatology) and the contribution of each component in our core architecture.
Subtracting climatology priors is critical for performance.
As mentioned in our methodology, we argue that focusing the model on learning ”anomalies” rather than the full state is crucial for capturing slow-changing dynamics. To verify this, we train a variant of NeuralOM without subtracting the climatological mean. The results, shown in Table 4, are striking. For sea salinity (S) across 23 depth layers, the model with climatology priors (W Subtract Climatology) achieved an ACC of 0.6412, whereas the model without them (W/O Subtract Climatology) only reached 0.1061—a performance gap of over 5-fold. This result provides strong evidence that incorporating physical priors like climatology into the input representation is a prerequisite for enabling the model to effectively learn fine-grained physical signals and achieve high-fidelity simulations.
Each core architectural component is indispensable.
Next, we evaluate the individual contributions of NeuralOM’s core components by removing them one by one. The results, summarized in Table 3, show that removing any single component leads to a significant degradation in performance:
Without Progressive Residual Correction (w/o PRC): When we replaced the multi-stage PRC framework with a single, deeper model, the RMSE increased from 0.6097 to 0.6692. This demonstrates that decomposing the complex prediction task into a series of simpler residual correction steps is crucial for suppressing error accumulation and ensuring long-term stability.
Without Physics-inspired Edge-Interactions (w/o PEI): Removing our physics-inspired interaction module (PEI), thereby reverting to a traditional, physics-agnostic message passing scheme, caused the ACC to drop from 0.5983 to 0.5168. This directly validates that explicitly modeling physical processes like gradients and couplings in message passing is effective for enhancing the model’s physical consistency.
Multi-scale Adaptive Node Aggregation (w/o MANA Sum/Mean): Removing either the sum or the mean aggregation path individually resulted in degraded performance. This confirms the necessity of our adaptive aggregation mechanism, which allows the model to flexibly handle ocean dynamics at different scales.
In summary, these ablation studies clearly demonstrate that the success of NeuralOM is not coincidental but is the result of a synergistic combination of its carefully designed components from the input representation and macroscopic framework to the microscopic interaction mechanisms.
Conclusion
In this paper, we introduce NeuralOM, a neural operator framework designed to solve the problem of long-term stability by simulating slow-changing physical systems. Its core innovations, a Progressive Residual Correction Framework and a Physics-Guided Graph Network, synergistically improve long-term stability and physical consistency. In the challenging task of global S2S ocean simulation, experiments demonstrate that NeuralOM significantly outperforms state-of-the-art models in forecast accuracy, long-term stability, and the ability to capture extreme events. Future work could explore integrating strict physical conservation laws into the loss function and reducing computational costs.
@@ A.1    Dataset
We conduct the experiments on GlORYS12 reanalysis data, offering daily mean data covering latitudes between -80° and 90° from 1993 to the present, which can be downloaded from https://data.marine.copernicus.eu. The subset we use includes years from 1993 to 2020, which is 1993-2017 for training, 2018-2019 for validating, and 2020 for testing. We use 4 depth level ocean variables (each with 23 depth levels, corresponding to 0.49 m, 2.65 m, 5.08 m, 7.93 m, 11.41 m, 15.81 m, 21.60 m, 29.44 m, 40.34 m, 55.76 m, 77.85 m, 92.32 m, 109.73 m, 130.67 m, 155.85 m, 186.13 m, 222.48 m, 266.04 m, 318.13 m, 380.21 m, 453.94 m, 541.09 m and 643.57 m), Sea salinity (S), Sea stream zonal velocity (U), Sea stream meridional velocity (V), Sea temperature (T), and 1 surface level variable Sea surface height (SSH). To improve computational efficiency, we use bilinear interpolation to downsampling them to 1/2 degree (H=361, W=720). And to better adapt to the input of different architecture models, we use the data with size 360 × 720. 4 atmosphere variables from ERA5 are used as forcing, which include 10 metre u wind component (U10M), 10 metre v wind component (V10M), 2 metre temperature (T2M), and mean sea level pressure (MSLP). The ERA5 data can be downloaded from https://cds.climate.copernicus.eu, the official website of Climate Data Store (CDS). All the data we used are shown in Table 5.
@@ A.2    Data Preprocessing
Table 5: The data details in this work.
B    Details of Physics-Guided Graph Network
B.1    Multi-scale Graph Encoder
We employ the multi-scale graph encoder to map the data from lat-lon grids into a graph structure. The multi-scale graph structure can be defined as:
$MLP=\mathcal{N}(\sigma(\mathcal{L}(\cdot))),$ (9)
where, $\mathrm{ESMLP}(\cdot)$ is the Edge Sum MLP (Pfaff et al. 2020). The mesh node features are updated by an MLP:
B.2    Physics-guided Adaptive Graph Messaging
To make the message passing more consistent with the evolution process of the ocean dynamics system, we propose a module called Physics-guided Adaptive Messaging (PAGM). PAGM includes a physics-inspired edge-interactions module and a multi-scale adaptive node aggregation module. In our setup, 16 PAGM are applied to conduct massaging.
Physics-Inspired Edge-Interactions.
In ocean simulation, ocean currents, temperature, salinity, and other factors interact with each other, and these interactions span different scales and layers. Unlike traditional MLP-based update strategy, which update edges by simply concatenating edges and nodes, we introduce a physics-inspired edge-interactions module to more accurately simulate the dynamic relationships in the ocean system. For the sender node feature $\mathbf{h}_{s(i)}$ and receiver nodes feature $\mathbf{h}_{r(i)}$, we first calculate their mutual interaction:
$\mathbf{h}_{d(i)}=\mathbf{h}_{s(i)}-\mathbf{h}_{r(i)},$ (14)
$\mathbf{h}_{mp(i)}=\mathbf{h}_{s(i)}*\mathbf{h}_{r(i)},$ (15)
$\mathbf{h}_{cos(i)}=cos(\mathbf{h}_{s(i)},\mathbf{h}_{r(i)}),$ (16)
$\mathbf{h}_{s(i)}=\mathcal{R}(\mathbf{h}_{s(i)},\mathbf{h}_{mp(i)},\mathbf{h}_{mp(i)},\mathbf{h}_{cos(i)}),$ (18)
$\mathbf{h}_{r(i)}=\mathcal{R}(\mathbf{h}_{r(i)},\mathbf{h}_{mp(i)},\mathbf{h}_{mp(i)},\mathbf{h}_{cos(i)}).$ (19)
${\mathcal{E}_{f}}={\mathcal{E}_{f}}^{\prime}+{\mathcal{E}_{f}}.$ (23)
Multi-scale Adaptive Node Aggregation.
Unlike traditional node update strategies, which typically sum or average the features of neighboring nodes, we consider the movements of the ocean at different scales. To achieve adaptive adjustment of aggregation methods across different scales, we introduce a multi-scale adaptive node aggregation module that dynamically adjusts the weights of the aggregation strategy. For rapid local ocean dynamics with smaller scales, the gating mechanism automatically tends to favor sum aggregation, capturing the local rapid dynamic changes quickly and accurately. Conversely, when the model identifies that the motion scale around the node is larger and the changes are slower, it tends to prefer mean aggregation, which achieves a smoother and more robust representation of the large-scale ocean state.
$h_{sum}^{\prime}=\mathrm{MLP}_{sum}(h,\sum\mathcal{E}_{f}^{\prime}).$ (24)
$h_{mean}^{\prime}=\mathrm{MLP}_{mean}(h,\sum\mathcal{E}_{f}^{\prime}).$ (25)
B.3    Decoder
$\mathcal{V}_{f}^{G}=\mathcal{V}_{f}^{G}+\mathcal{V}_{f}^{G^{\prime}}.$ (30)
Finally, we use an MLP to predict the next step results:
$\hat{O}_{t+1}=\operatorname{MLP}_{d2}\left(\mathcal{V}_{f}^{G}\right).$ (31)
C    Experiments Details
C.1    Evaluation Metrics
where $\mathbf{A}_{i,j,t}^{\mathcal{K}}$ represents the value of variable $\mathcal{J}$ at horizontal coordinate $(i,j)$ and time t. Latitude-dependent weights are defined as $L(i)=N_{\text{lat }}\times\frac{\cos\phi_{i}}{\sum_{i’=1}^{N_{\text{lat}}}\cos\phi_{i’}}$, where $\phi_{i}$ is the latitude at index i. The anomaly of $A$, denoted as $A^{\prime}$, is computed as the deviation from its climatology, which corresponds to the long-term mean of the meteorological state estimated from multiple years of training data. RMSE and ACC are averaged across all time steps and spatial coordinates, providing summary statistics for each variable $\mathcal{J}$ at a given lead time $\Delta t$.
where, true positives (TP) represent the number of instances in which the ocean state is correctly simulated, while false positives (FP) and false negatives (FN) are defined analogously. $F=\frac{\mathrm{FP}}{\mathrm{FP}+\mathrm{TP}}$ is the false alarm rate, and $H=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}}$ represents the hit rate. In the simulation task, we report the average results for 240 initial conditions (ICs), and in forecasting, for 50 ICs.
C.2    Model Training